# NuGraph2 with Context–Aware Inputs: Physics–Inspired Improvements in Semantic Segmentation

### v1

@@ 1 Introduction
Graph neural networks (GNNs) have emerged as powerful tools in particle physics for analyzing detector data, largely because they can natively represent physical relationships in irregular data [1, 2, 3]. Unlike traditional image or sequence models, GNNs allow high-level physical features to be attached to measured data and analyzed by the model [4]. Recent work has leveraged this flexibility to inject domain knowledge at various stages of the learning pipeline [5]. Among these approaches, three pertinent strategies are: enriching inputs with physics-inspired features, multi-task learning with auxiliary outputs, and physics-informed loss constraints. These approaches have demonstrated improved expressivity, disambiguation of classes in sparse data, and physically plausible predictions in tasks relevant to Liquid Argon Time Projection Chambers (LArTPC) detectors and beyond, providing key inspiration for our method.
@@
One line of research augments the raw graph inputs with engineered features that encode geometric or physical context. By summarizing local structures or known physics, these descriptors can enhance GNN expressivity, especially when data are sparse or ambiguous. For example, Drielsma et al. [6] introduce 22 geometric descriptors per node (e.g., a fragment’s covariance, principal axes, centroid, etc.) in their Graph Particle Aggregator (GrapPA) to help assemble electromagnetic shower fragments in LArTPC data, where a fragment is fully connected cluster of energy deposits (hits) in the detector. Pairwise relationships are likewise encoded: each edge carries 19 physics-motivated features (such as closest approach distances between fragments) to inform the network which pieces likely belong to the same particle shower. These hand-crafted features provide the network with a notion of shape, orientation, and proximity that would be difficult to learn from raw data alone, leading to significantly improved clustering of particle showers.
@@
More generally, incorporating domain-specific inputs has proven beneficial in other particle physics ML applications. For instance, by combining 3D convolutional neural networks with physics-inspired statistical summaries of the electromagnetic radiation pattern, Kiesler et al. [7] demonstrate that incorporating domain knowledge significantly enhances the precision of muon energy estimates. Their results highlight the complementary benefits of deep learning and hand-crafted, physics-driven features in this context. Overall, encoding physically relevant attributes at the input stage guide GNNs to learn representations that respect known structures in the data (e.g., detector geometry or particle shower morphology). This idea motivates our inclusion of domain-inspired graph features to provide structural context beyond the raw hit information.
@@
Another strategy is to design networks that jointly predict multiple targets to exploit known correlations among classes or tasks. In general, multi-task learning has been found to improve performance in physics machine learning by providing internal consistency checks: the network is pressured to find representations useful for several related predictions rather than overfit to one. In high-energy physics, many labels are interdependent (for example, the presence of certain particle types in an event is correlated with the event’s topology or origin). Rather than training separate models, a single network with auxiliary decoders for related tasks can share representations and teach the model these inter-class relationships. A notable example is the DUNE Collaboration’s Convolutional Visual Network (CVN), which uses a Convolutional Neural Network (CNN) to simultaneously classify neutrino flavor, interaction type, and particle content of each event [8]. The CVN produces multiple sets of output probabilities for neutrino flavor (electron, muon, tau, or none), for interaction channel (e.g., charged-current vs. neutral-current), and even for the counts of protons, pions, and neutrons in the final state. This multi-head architecture leverages the fact that, for instance, an electron-neutrino charged-current event is likely to contain an outgoing electron and proton, whereas a neutral-current event might produce different secondaries. By learning all these labels together, the network’s shared hidden features capture the semantic relationships between flavor and topology, yielding more robust classification than any single-task model.
@@
Beyond model architecture, an important avenue is incorporating physics domain knowledge into the loss function or training objective. By adding custom loss terms or regularization penalties, one can directly encourage the model to produce physically plausible outputs or to respect known laws [9, 10, 11]. In other words, the idea is to use the loss function as a steering mechanism, gently constraining the model toward physically consistent behavior without hard-coding it. This concept underlies the growing field of physics-informed neural networks (PINNs), where differential equations or other theoretical constraints are imposed as part of the loss [12]. By design, such constraints drive the learned model towards the manifold of solutions allowed by fundamental physics, improving generalization and trustworthiness of the predictions [13]. Even when exact laws are not known or easily applied, domain-informed loss functions can be crafted to encode experiment-specific expectations. In the context of graph networks, Sharma et al. [14] recently demonstrated a physics-informed GNN for multi-body dynamics that explicitly enforces conservation laws. Their model adds terms to the loss to conserve linear and angular momentum in sphere collisions, yielding predictions that obey symmetries and remain stable over long simulations.
@@
Drawing inspiration from these prior works, our study investigates the impact of injecting physics-domain knowledge into the NuGraph2 [15] network through tailored input features, auxiliary decoders, and loss regularization terms. In a LArTPC, three wire planes ($u$, $v$, $y$), arranged at different wire directions, collect ionization electrons drifting through the liquid argon. In NuGraph2, each detector hit is represented as a node. Delaunay triangulation algorithm is used to connect nodes belonging to the same plane (planar nodes), while a set of nexus nodes serves as inter-plane connection points, enabling information exchange between the different wire planes. NuGraph2 labels each detector hit (i.e. a planar node) based on a binary discriminator for signal vs background and a semantic classifier, where classes correspond to five different particle types: Minimum Ionizing Particles (MIPs), Highly Ionizing Particles (HIPs), electro-magnetic showers, Michel electrons, and diffuse activity. A more detailed introduction to NuGraph2 is in [16].
Our primary goal is to improve the network’s ability to correctly label Michel electron hits, whose classification performance is significantly lower than that of other classes due to their underrepresentation in the dataset. While our main focus is on Michel electrons, we also examine how Michel-targeted interventions may affect the classification accuracy of other particle classes. Similar to the NuGraph2 result [15], our work is also based on the MicroBooNE public data sets [17, 18], which include simulated neutrino interactions from the Booster Neutrino Beam on top of off-beam data collected with the MicroBooNE detector [19]. More information on these data sets can be found in [20]. The code implementation is available on Github at https://github.com/vitorgrizzi/nugraph_phys/tree/main_phys.
@@ 2.1 Feature Extension
Recent work has revealed that standard GNNs, particularly those based on message passing, are inherently limited in their ability to capture graph structure. In particular, they are proved to be no more expressive than the 1-Weisfeiler-Leman graph isomorphism test and consequently fail to distinguish graph substructures or encode higher-order relational context among nodes [21]. This means that GNNs are generally unaware of the structural roles of nodes and cannot capture meaningful patterns such as cycles, cliques, or neighborhood asymmetries, which are often critical in physics-based or scientific domains.
To mitigate the limitations of GNNs in learning contextual information from both the structural roles of nodes and the topological properties of the graph, we engineered features that encode each node’s relational and structural context within the graph. These features, derived from domain knowledge, include quantities such as node degree, local geometric distances, and neighbor-relative differences in wirevstimewire\ vs\ time space. By doing so, we aim to inject inductive biases that GNNs would otherwise struggle to learn from data alone, thereby enhancing the model’s expressivity and its ability to resolve class ambiguities, especially in challenging cases such as Michel electron identification.
@@
  • Shortest edge length in the wirevstimewire\ vs\ time plane.
  • Double difference between the wirewire value of the current node and its two nearest neighbors in the wirevstimewire\ vs\ time plane. This feature is denoted by $\Delta wire$.
  • Double difference between the timetime value of the current node and its two nearest neighbors in the wirevstimewire\ vs\ time plane. This feature is denoted by $\Delta time$.
The node degree, defined as the number of edges connected to a node, explicitly encodes structural information about that node’s role in the graph. It helps identify “hub“ nodes [16] that connect to many others, thereby influencing a broader neighborhood. Conversely, nodes with lower degree tend to exchange information with fewer neighbors and thus represent more localized and specific features. An example of the graph connectivity is shown in Figure 1. The rationale for including the node degree as an extra feature is that it may help the network modulate the contribution of a node during the message passing phase based on the connectedness or specificity of a node.
Figure 1: Node connectivity in the wirevstimewire\ vs\ time plane for the three different planes $u$, $v$, and $y$
The second feature is the Euclidean distance to the nearest neighbor in the wirevstimewire\ vs\ time plane, which helps the network distinguish between isolated hits and those that are part of continuous tracks. This “isolation degree" is particularly relevant for semantic classification and filtering tasks, as expected from physics intuition. For example, minimum ionizing particles (MIPs) and highly ionizing particles (HIPs) form coherent tracks in the wirevstimewire\ vs\ time plane; thus, isolated hits are unlikely to belong to either category.
The $\Delta wire$ and $\Delta time$ features also assist in track identification. For each node $i$, the difference is calculated as
@@
where $x_{i}$ is the coordinate (either wirewire or timetime) of the reference node, and $x_{j1}$ and $x_{j2}$ are the coordinates of its two nearest neighbors in the wirevstimewire\ vs\ time plane, using Euclidean distance. The double differences $\Delta wire$ and $\Delta time$ remain approximately zero along a track (or any segment with fixed direction). Therefore, this metric can help the network determine whether a node is part of a track, which in turn constrains the likelihood of it belonging to a specific class.
@@
Figure 3 demonstrates the classification results obtained by our physics-based extended features. It can be seen that both precision and recall metrics are systematically improved compared to the baseline model’s results shown in Figure 2. The largest gain is observed for Michel electrons, which initially had the worst performance due to their under-representation in the dataset. The additional information carried by the new features enabled the network to better extract relevant signals from the few events containing Michel electrons, thereby improving their recognition. An alternative explanation stems from the observation that Michel electrons occupy regions of latent space that overlap with other classes, particularly MIPs. The introduction of additional features provides the semantic head with auxiliary information that helps to better disentangle these overlapping regions in the latent representation space. Overall, this positive result highlights that incorporating context-aware features, crafted using physics intuition, is an effective strategy for injecting domain knowledge and enhancing the network’s performance.
@@
While the addition of these features provide improved performance, it may bring an increased risk of domain shift between the training and application datasets. In fact, application datasets which match the training in terms of the nominal (local) input node features, may instead show a discrepancy in terms of the context information added by the new features. For instance, the minimum Euclidean distance to the nearest neighbor may carry information about the global direction of the particle producing the nodes, projected to the specific plane where the nodes are located. The MicroBooNE detector [19] is exposed to two different neutrino beams entering the detector from different directions. Each beam thus has a preferential direction at which charged particles observed in the detector are produced, i.e. along the neutrino direction. The usage of this feature can thus introduce a domain shift when training on a dataset from one beam and applied to a dataset from the other beam. We also note that the definition of the $\Delta wire$ and $\Delta time$ features is designed to minimize this type of domain shift: in fact the double difference in Eq. 1 will highlight differences in direction at the node, while a definition based on a single difference, e.g. $\Delta x=x_{j1}-x_{j2}$, would carry information that relates more directly to the global particle direction and may cause domain shift problems.
@@ 2.3 Michel Energy Regularization
Michel electrons are the product of the decay of muons at rest in the detector. Their energy spectrum is governed by energy conservation laws applied to the kinematics of the three-body decay process [22]. To improve the Michel electron metrics, we introduced a physics-based regularization term in the loss function that penalizes the model for classifying nodes as Michel electrons when the sum of the energies deposited by these nodes exceeds a threshold or deviates from the expected energy distribution observed in the batch. The key idea is that while a single Michel electron can produce a variable number of detector hits in each LArTPC plane, the total deposited energy across those hits is ultimately constrained by the energy of the originating Michel electron [23]. We tested three different regularization schemes:
@@
The first scheme penalizes the network for assigning the Michel label to nodes whose deposited energy exceeds the expected range for Michel electrons, as observed in the dataset. Since hits in our dataset do not measure the deposited energy directly, we used one of their features as a proxy, i.e. the waveform integral ($Integral$). Figure 4 shows the distribution of Michel electron hit integral vs. deposited energy; as expected, they are approximately linearly related, with the slope of the fitted black line encoding this relationship. This allowed us to estimate the deposited energy from the $Integral$ hit feature and apply the regularization. This regularization is expressed as a penalty term to the loss function:
@@
We also explored an alternative regularization scheme based on how well the Michel energy distribution matched theoretical and simulation-based expectations. Figure 5 shows the expected deposited energy spectrum of Michel electrons, summing contributions across all LArTPC planes. The peak near $E=0$ MeV is an artifact of the simulation related to detector boundary effects and unresponsive regions, not a fundamental physics feature. Accordingly, the Langauss fit (red curve) was applied only for $E\geq 8.5$ MeV. We penalized the model for deviations between the Michel energy distribution and either the Langauss fit or the empirical simulation distribution.
@@
Two other limitations in this approach that likely contribute to the degradation in performance are the following. The first is related to the relationship between the $Integral$ feature and deposited energy. Although Figure 4 shows a clear positive linear trend for Michel hits, the large spread of data points around the fitted line indicates significant variance. This implies that while $Integral$ and energy are correlated, the uncertainty in this relationship limits the reliability of energy estimation based solely on $Integral$, thereby undermining the effectiveness of energy-based regularization. The second limitation relies in the fact that NuGraph2 only makes predictions about hits, and does not have a concept of a particle instance with a given deposited energy. Consequently, there is not a one-to-one correspondence between a set of nodes semantically labeled according to a given class and a specific particle unless there is only a single particle of that class in the event. This means that the penalty term is computed for each event in the batch, which represents a coarser granularity than a more natural per-particle contribution. Improving these limitation is beyond the scope of this paper, as it will require completion of ongoing updates to NuGraph that will allow to predict per-particle attributes.

### v3

@@ 1 Introduction
Graph neural networks (GNNs) have emerged as powerful tools in particle physics for analyzing detector data, largely because they can natively represent physical relationships in irregular data [bronstein2017geometric, cao2020comprehensive, atz2021geometric]. Unlike traditional image or sequence models, GNNs allow high-level physical features to be attached to measured data and analyzed by the model [thais2022graph]. Recent work has leveraged this flexibility to inject domain knowledge at various stages of the learning pipeline [dash2022review]. Among these approaches, three pertinent strategies are: enriching inputs with physics-inspired features, multi-task learning with auxiliary outputs, and physics-informed loss constraints. These approaches have demonstrated improved expressivity, disambiguation of classes in sparse data, and physically plausible predictions in tasks relevant to Liquid Argon Time Projection Chambers (LArTPC) detectors and beyond, providing key inspiration for our method.
@@
One line of research augments the raw graph inputs with engineered features that encode geometric or physical context. By summarizing local structures or known physics, these descriptors can enhance GNN expressivity, especially when data are sparse or ambiguous. For example, Drielsma et al. [drielsma2021clustering] introduce 22 geometric descriptors per node (e.g., a fragment’s covariance, principal axes, centroid, etc.) in their Graph Particle Aggregator (GrapPA) to help assemble electromagnetic shower fragments in LArTPC data, where a fragment is fully connected cluster of energy deposits (hits) in the detector. Pairwise relationships are likewise encoded: each edge carries 19 physics-motivated features (such as closest approach distances between fragments) to inform the network which pieces likely belong to the same particle shower. These hand-crafted features provide the network with a notion of shape, orientation, and proximity that would be difficult to learn from raw data alone, leading to significantly improved clustering of particle showers.
@@
More generally, incorporating domain-specific inputs has proven beneficial in other particle physics ML applications. For instance, by combining 3D convolutional neural networks with physics-inspired statistical summaries of the electromagnetic radiation pattern, Kiesler et al. [kieseler2022calorimetric] demonstrate that incorporating domain knowledge significantly enhances the precision of muon energy estimates. Their results highlight the complementary benefits of deep learning and hand-crafted, physics-driven features in this context. Overall, encoding physically relevant attributes at the input stage guide GNNs to learn representations that respect known structures in the data (e.g., detector geometry or particle shower morphology). This idea motivates our inclusion of domain-inspired graph features to provide structural context beyond the raw hit information.
@@
Another strategy is to design networks that jointly predict multiple targets to exploit known correlations among classes or tasks. In general, multi-task learning has been found to improve performance in physics machine learning by providing internal consistency checks: the network is pressured to find representations useful for several related predictions rather than overfit to one. In high-energy physics, many labels are interdependent (for example, the presence of certain particle types in an event is correlated with the event’s topology or origin). Rather than training separate models, a single network with auxiliary decoders for related tasks can share representations and teach the model these inter-class relationships. A notable example is the DUNE Collaboration’s Convolutional Visual Network (CVN), which uses a Convolutional Neural Network (CNN) to simultaneously classify neutrino flavor, interaction type, and particle content of each event [abi2020neutrino]. The CVN produces multiple sets of output probabilities for neutrino flavor (electron, muon, tau, or none), for interaction channel (e.g., charged-current vs. neutral-current), and even for the counts of protons, pions, and neutrons in the final state. This multi-head architecture leverages the fact that, for instance, an electron-neutrino charged-current event is likely to contain an outgoing electron and proton, whereas a neutral-current event might produce different secondaries. By learning all these labels together, the network’s shared hidden features capture the semantic relationships between flavor and topology, yielding more robust classification than any single-task model.
@@
Beyond model architecture, an important avenue is incorporating physics domain knowledge into the loss function or training objective. By adding custom loss terms or regularization penalties, one can directly encourage the model to produce physically plausible outputs or to respect known laws [nabian2020physics, nabian2021efficient, djeumou2022neural]. In other words, the idea is to use the loss function as a steering mechanism, gently constraining the model toward physically consistent behavior without hard-coding it. This concept underlies the growing field of physics-informed neural networks (PINNs), where differential equations or other theoretical constraints are imposed as part of the loss [raissi2019physics]. By design, such constraints drive the learned model towards the manifold of solutions allowed by fundamental physics, improving generalization and trustworthiness of the predictions [thangamuthu2022unravelling]. Even when exact laws are not known or easily applied, domain-informed loss functions can be crafted to encode experiment-specific expectations. In the context of graph networks, Sharma et al. [sharma2025dynami] recently demonstrated a physics-informed GNN for multi-body dynamics that explicitly enforces conservation laws. Their model adds terms to the loss to conserve linear and angular momentum in sphere collisions, yielding predictions that obey symmetries and remain stable over long simulations.
@@
Drawing inspiration from these prior works, our study investigates the impact of injecting physics-domain knowledge into the NuGraph2 [nugraph2] network through tailored input features, auxiliary decoders, and loss regularization terms. In a LArTPC, three wire planes ($u$, $v$, $y$), arranged at different wire directions, collect ionization electrons drifting through the liquid argon. In NuGraph2, each detector hit is represented as a node. Delaunay triangulation algorithm is used to connect nodes belonging to the same plane (planar nodes), while a set of nexus nodes serves as inter-plane connection points, enabling information exchange between the different wire planes. NuGraph2 labels each detector hit (i.e. a planar node) based on a binary discriminator for signal vs background and a semantic classifier, where classes correspond to five different particle types: Minimum Ionizing Particles (MIPs), Highly Ionizing Particles (HIPs), electro-magnetic showers, Michel electrons, and diffuse activity. A more detailed introduction to NuGraph2 is in [explainability].
[…]
@@ 2.1 Feature Extension
Recent work has revealed that standard GNNs, particularly those based on message passing, are inherently limited in their ability to capture graph structure. In particular, they are proved to be no more expressive than the 1-Weisfeiler-Leman graph isomorphism test and consequently fail to distinguish graph substructures or encode higher-order relational context among nodes [bouritsas2022improving]. This means that GNNs are generally unaware of the structural roles of nodes and cannot capture meaningful patterns such as cycles, cliques, or neighborhood asymmetries, which are often critical in physics-based or scientific domains.
To mitigate the limitations of GNNs in learning contextual information from both the structural roles of nodes and the topological properties of the graph, we engineered features that encode each node’s relational and structural context within the graph. These features, derived from domain knowledge, include quantities such as node degree, local geometric distances, and neighbor-relative differences in timevswiretime\ vs\ wire space. By doing so, we aim to inject inductive biases that GNNs would otherwise struggle to learn from data alone, thereby enhancing the model’s expressivity and its ability to resolve class ambiguities, especially in challenging cases such as Michel electron identification.
@@
  • Shortest edge length in the timevswiretime\ vs\ wire plane.
  • Double difference between the wirewire value of the current node and its two nearest neighbors in the timevswiretime\ vs\ wire plane. This feature is denoted by $\Delta wire$.
  • Double difference between the timetime value of the current node and its two nearest neighbors in the timevswiretime\ vs\ wire plane. This feature is denoted by $\Delta time$.
The inspiration to include these features was guided by inspecting the neutrino-interaction event graphs. Figure 1 illustrates one such event in the timevswiretime\ vs\ wire plane for the $u$, $v$, and $y$ detector planes. It shows coherent linear tracks, characteristic of HIPs and MIPs, and nodes with very distinct structural roles in the graph, as indicated by their connectivity and nearest-neighbor distance. These patterns were key to the construction of the additional input features.
The node degree, defined as the number of edges connected to a node, explicitly encodes structural information about that node’s role in the graph. It helps identify “hub“ nodes [explainability] that connect to many others and therefore influence a broader neighborhood. Conversely, nodes with lower degree tend to exchange information with fewer neighbors and thus represent more localized and specific features. Thus, including the node degree as an extra feature may help the network modulate the contribution of a node during the message passing phase based on its connectedness or specificity.
Figure 1: Node connectivity in the timevswiretime\ vs\ wire plane for the three different planes $u$, $v$, and $y$. Each node represents a detector hit and edges were constructed using Delaunay triangulation algorithm. Nodes are color-coded according to their true semantic label. The graph exhibits linear tracks characteristic of HIPs and MIPs, and dispersed nodes with high connectivity acting as “hubs“. Both patterns served as inspiration for augmenting the input features.
The second feature is the Euclidean distance to the nearest neighbor in the timevswiretime\ vs\ wire plane, which helps the network distinguish between isolated hits and those that are part of continuous tracks. This “isolation degree" is particularly relevant for semantic classification and filtering tasks, as expected from physics intuition. For example, minimum ionizing particles (MIPs) and highly ionizing particles (HIPs) form coherent tracks in the timevswiretime\ vs\ wire plane; thus, isolated hits are unlikely to belong to either category.
Finally, the $\Delta wire$ and $\Delta time$ features also assist in track identification. For each node $i$, the difference is calculated as
@@
where $x_{i}$ is the coordinate (either wirewire or timetime) of the reference node, and $x_{j1}$ and $x_{j2}$ are the coordinates of its two nearest neighbors in the timevswiretime\ vs\ wire plane, using Euclidean distance. The double differences $\Delta wire$ and $\Delta time$ remain approximately zero along a track (or any segment with fixed direction) as one can see from Figure 1. Therefore, this metric can help the network determine whether a node is part of a track, which in turn constrains the likelihood of it belonging to a specific class.
@@
Figure 3 demonstrates the classification results obtained by our physics-based extended features as Precision and Recall matrices. Precision is defined as $TP/(TP+FP)$ where $TP$ refers to True Positive and $FP$ to False Positive, and indicates the percentage of particles assigned to class $K$ that actually belongs to $K$. On the other hand, recall is defined as $TP/(TP+FN)$ where $FN$ refers to False Negative, and indicates the percentage of particles from class $K$ that were correctly labeled as $K$. These metrics can also be referred to as purity and efficiency, respectively. Nevertheless, Figure 3 shows that both precision and recall metrics are systematically improved when compared to the baseline model’s results shown in Figure 2. The largest gain is observed for Michel electrons, which initially had the worst performance due to their underrepresentation in the dataset. The additional information carried by the new features enabled the network to better extract relevant signals from the few events containing Michel electrons, thereby improving their recognition. An alternative explanation stems from the observation that Michel electrons occupy regions of latent space that overlap with other classes, particularly MIPs. The introduction of additional features provides the semantic head with auxiliary information that helps to better disentangle these overlapping regions in the latent representation space. Overall, this positive result highlights that incorporating context-aware features, crafted using physics intuition, is an effective strategy for injecting domain knowledge and enhancing the network’s performance.
@@
While the addition of these features provide improved performance, it may bring an increased risk of domain shift between the training and application datasets. In fact, application datasets which match the training in terms of the nominal (local) input node features, may instead show a discrepancy in terms of the context information added by the new features. For instance, the minimum Euclidean distance to the nearest neighbor may carry information about the global direction of the particle producing the nodes, projected to the specific plane where the nodes are located. The MicroBooNE detector [acciarri2017design] is exposed to two different neutrino beams entering the detector from different directions. Each beam thus has a preferential direction at which charged particles observed in the detector are produced, i.e. along the neutrino direction. The usage of this feature can thus introduce a domain shift when training on a dataset from one beam and applied to a dataset from the other beam. We also note that the definition of the $\Delta wire$ and $\Delta time$ features is designed to minimize this type of domain shift: in fact the double difference in Eq. 1 will highlight differences in direction at the node, while a definition based on a single difference, e.g. $\Delta x=x_{j1}-x_{j2}$, would carry information that relates more directly to the global particle direction and may cause domain shift problems.
@@ 2.3 Michel Energy Regularization
Michel electrons are the product of the decay of muons at rest in the detector. Their energy spectrum is governed by energy conservation laws applied to the kinematics of the three-body decay process [ParticleDataGroup:2024cfk]. To improve the Michel electron metrics, we introduced a physics-based regularization term in the loss function that penalizes the model for classifying nodes as Michel electrons when the sum of the energies deposited by these nodes exceeds a threshold or deviates from the expected energy distribution observed in the batch. The key idea is that while a single Michel electron can produce a variable number of detector hits in each LArTPC plane, the total deposited energy across those hits is ultimately constrained by the energy of the originating Michel electron [MicroBooNE:2017kvv]. We tested three different regularization schemes:
@@
The first scheme penalizes the network for assigning the Michel label to nodes whose deposited energy exceeds the expected range for Michel electrons, as observed in the dataset. Since hits in our dataset do not measure the deposited energy directly, we used one of their features as a proxy, i.e. the waveform integral ($Integral$). This correlation is modeled through a chain of different tools. The propagation of particles in the detector volume is simulated with Geant4 [GEANT4:2002zbu] v10, which produces a collection of true energy deposits in the detector volume. The corresponding ionization charge is then propagated in the detector using various tools available in LArSoft [Snider:2017wjd]. The recombination of electron-ion pairs is modeled with a modified box model [ArgoNeuT:2013kpa]. Distortions in the electric field due to the space-charge effects are simulated with data-driven electric field maps [MicroBooNE:2019koz, MicroBooNE:2020kca]. Other non-uniformities, resulting from variations in electron lifetime or wire response are also simulated. The readout electronics and wire response are modeled through a simulation of the charge induced by electrons that drift to the wire planes at the anode [MicroBooNE:2018swd, MicroBooNE:2018vro]. Waveforms are then processed through a reconstruction chain that includes noise filtering [MicroBooNE:2017qiu], signal deconvolution and region-of-interest finding [MicroBooNE:2018swd]. Hits [Baller:2017ugz, Berkman:2021ffy] are Gaussian fits to peaks in regions of interest on the deconvolved waveform. Their Integral, expressed in ADC counts, is computed from the corresponding best-fit Gaussian function parameters. In addition to the instrumental effects listed above, deviations from an ideal linear dependence can also be due to other contributions to the hit integral in addition to the true Michel energy deposition. In particular, other particles, such as cosmics, with trajectories overlapping with Michel hits in the wire readout plane can lead to a substantial bias towards larger integral values.
Figure 4 shows the distribution of Michel electron hit integral vs. deposited energy; as expected, they are approximately linearly related, with the slope of the fitted black line encoding this relationship. This allowed us to estimate the deposited energy from the $Integral$ hit feature and apply the regularization. This regularization is expressed as a penalty term to the loss function:
@@
We also explored an alternative regularization scheme based on how well the observed Michel energy distribution matches predictions from our simulation. Figure 5 shows the expected deposited energy spectrum of Michel electrons, summing contributions across all LArTPC planes. The peak near $E=0$ MeV is due to Michel electrons propagating mostly outside the detector fiducial volume or in regions dominated by unresponsive wires. We use two methods to model the expected Michel energy. In the first method, we aim at modeling the probability based on a functional form that accounts for cases where the Michel deposits most of its energy in the detector. We therefore fit the histogram in Figure 5 with a Langauss fit (red curve), computed only for $E\geq 8.5$ MeV. In contrast, the second method uses the actual histogram in Figure 5 to define a probability density function that includes all detector effects, including the peak at 0 MeV. We penalized the model for differences between the Michel energy distribution obtained from the NuGraph predictions and the expected distribution from the simulation, computed either with the Langauss fit or using the empirical simulation distribution.
@@
Two other limitations in this approach that likely contribute to the degradation in performance are the following. The first is related to the relationship between the $Integral$ feature and deposited energy. Although Figure 4 shows a clear positive linear trend for Michel hits, the large spread of data points around the fitted line indicates significant variance. This implies that while $Integral$ and energy are correlated, the uncertainty in this relationship limits the reliability of energy estimation based solely on $Integral$, thereby undermining the effectiveness of energy-based regularization. A similar behavior is observed in Figure 5, where the Langauss fit, represented by the red line, fails to accurately capture the fast decay and extended tail of the data distribution above the 80 MeV region. However, note that most of the probability mass is concentrated below that region. Thus, although the fit is less accurate in the region beyond 80 MeV, it is still adequate for a proof-of-principle, and the residual mismatch has little influence on our conclusions. The second limitation relies in the fact that NuGraph2 only makes predictions about hits, and does not have a concept of a particle instance with a given deposited energy. Consequently, there is not a one-to-one correspondence between a set of nodes semantically labeled according to a given class and a specific particle unless there is only a single particle of that class in the event. This means that the penalty term is computed for each event in the batch, which represents a coarser granularity than a more natural per-particle contribution. Improving these limitation is beyond the scope of this paper, as it will require completion of ongoing updates to NuGraph that will allow to predict per-particle attributes.